# Assessing Vulnerability in Smart Contracts: The Role of Code Complexity Metrics in Security Analysis

### v3

@@ Abstract
Codes with specific characteristics are more exposed to security vulnerabilities. Studies have revealed that codes that do not stick to best practices are more challenging to verify and maintain, increasing the likelihood of unnoticed or unintentionally introduced vulnerabilities. Given the crucial role of smart contracts in blockchain systems, ensuring their security and conducting thorough vulnerability analysis is critical. This study investigates the use of code complexity metrics as indicators of vulnerable code in Solidity smart contracts. We highlight the significance of complexity metrics as valuable complementary features for vulnerability assessment and provide insights into the individual power of each metric. By analyzing 21 complexity metrics, we explored their interrelation, association with vulnerability, discriminative power, and mean values in vulnerable versus neutral codes. The results revealed some high correlations and potential redundancies among certain metrics, but weak correlations between each independent metric and vulnerability. Nevertheless, we found that all metrics can effectively discriminate between vulnerable and neutral codes, and most complexity metrics, except for three, exhibited higher values in vulnerable codes.
Index Terms:
Smart contract, Software Metrics, Complexity Metrics, Vulnerability detection, Statistical Analysis
@@ I Introduction & Motivation
Objectivity and Quantifiability Complexity metrics provide an objective and reproducible way to measure code characteristics [18]. Unlike qualitative assessments (e.g., code readability or subjective security evaluations), these metrics can be systematically computed, analyzed, and compared across large datasets. This makes them a suitable choice for large-scale vulnerability detection efforts.
Practical Integration into Automated Tools Automated vulnerability detection tools rely on measurable features to assess risk [19]. Complexity metrics are easy to extract from code and can be directly integrated into static analysis tools.
@@
smart contract Loss(\$)
UwU Lend 19.3M
Indodax 22M
BinX Exchange 26M
Penpie 27M
Hedgey Finance 44M
BtcTurk Hot Wallet 55M
DMM Bitcoin Private Key 125M
Orbit Chain Bridge 197M
WazirX 230M
PlayDapp 290M
Total 1.03B
@@ II Contribution & Research Questions
  • RQ4: How different are the metrics in vulnerable codes compared to neutral ones? We intend to examine whether higher complexity in code is linked to a higher likelihood of vulnerability occurrence. This approach enables a more targeted and efficient allocation of resources toward securing smart contracts.
@@ III-C Smart Contract Architecture
IV Complexity Metrics
@@ VI Experimental Setup
Paired T-test: A paired t-test, also known as a dependent t-test or a matched-pairs t-test, is a statistical hypothesis test used to compare the means of two related groups [42]. It is designed for situations where you have two sets of measurements on the same subjects or items, and the values in each set are paired together. The pairing can occur because the measurements are taken from the same individual or item under different conditions or at different time points.
The paired t-test is used to determine if there is a significant difference between the means of the two sets of paired data. The assumption behind this test is that the differences between the paired data points are normally distributed.
In this test, If the p-value is smaller than the chosen significance level (in our case 0.05), you reject the null hypothesis and conclude that there is a significant difference between the means of the two paired groups [42].
@@
VII Experimental Results
In this results section, we give answers to each research question asked. For a full report, you can look at our implementation’s source code which is publicly available at Zenodo [^4] https://zenodo.org/records/13364187 and GitHub [^5] https://github.com/MasoudJTehrani/SC-Complexity-Vulnerability
  • NOI and Avg. NOI: Are equal due to the limited use of invocations in smart contracts, suggesting that only one metric might be necessary. The high correlation between NOI and other metrics of SLOC, NF, WMC, and NOS was mentioned previously.
TABLE IV: Correlations between complexity metrics and existence of Vulnerabilities
@@ VII Experimental Results
Metrics SC Corr TP Corr
@@
RQ3: Are the metrics different in vulnerable and neutral codes? We sought to investigate whether the complexity metrics can effectively distinguish between functions with reported vulnerabilities and those without reported vulnerabilities. To assess whether the metrics in each group (vulnerable and neutral) can effectively discriminate between vulnerable and neutral codes, we employed the paired t-test. This statistical test was conducted by comparing the metric values of the vulnerable group with their corresponding values in the neutral group. A p-value greater than 0.05 indicates that the null hypothesis is true, suggesting that the metric’s values are derived from the same distribution and cannot effectively discriminate between vulnerable and neutral codes. On the other hand, a p-value less than or equal to 0.05 suggests that there are significant differences in the metric values between the two groups, indicating the potential of the metric to discriminate vulnerable codes from neutral codes. The p-values and the results are available in Table VI
TABLE VI: The paired t-test between each metric’s values in vulnerable and its corresponding values in neutral codes
Metrics Discriminative P-Value
SLOC yes <0.05
LLOC yes <0.05
CLOC yes <0.05
NF yes <0.05
WMC yes <0.05
NL yes <0.05
NLE yes <0.05
NUMPAR yes <0.05
NOS yes <0.05
DIT yes <0.05
NOA yes <0.05
NOD yes <0.05
CBO yes <0.05
NA yes <0.05
NOI yes <0.05
Avg. McCC yes <0.05
Avg. NL yes <0.05
Avg. NLE yes <0.05
Avg. NUMPAR yes <0.05
Avg. NOS yes <0.05
Avg. NOI yes <0.05
The table shows that all metrics can be sufficiently discriminative in distinguishing vulnerable codes from neutral codes in Solidity smart contracts.
@@
To answer this question, we need to compare the mean of each group. To compare the means of each group, we employed confidence intervals, ensuring that the resulting intervals encompass the true population parameter in around 95% of cases. This statistical approach allowed us to effectively evaluate the differences between vulnerable and neutral codes in Solidity smart contracts. The comparison results are presented in Figure 5.
@@ VIII-C Future Work
In future work, we have identified several areas for improvement and expansion in our study. One of our objectives is to include additional complexity metrics that are specific to smart contracts, such as the amount of gas expended, to enhance our analysis.
@@
Furthermore, we plan to explore the inclusion of comprehension complexity metrics, such as cognitive complexity [43, 44], in our evaluation. Cognitive complexity measures the difficulty of understanding code by considering factors such as nested conditions, control flow, and human cognition. By including such metrics, we can gain a better understanding of the cognitive load imposed on developers and its application in vulnerabilities of Smart Contracts. These future efforts will contribute to a more comprehensive and robust assessment of Smart Contract security.
@@ IX-B Internal Threats
This paper employed statistical hypothesis tests, including paired t-tests, to identify significant differences between paired data sets. Since these tests rely on several assumptions, their reliability—particularly for RQ3—may be affected if those assumptions are not met in a different setup. To mitigate this, these methods were selected based on prior research and the characteristics of the dataset. This also ensured the consistency with studies used for comparison.
@@ X Conclusion
Throughout our investigation, we have obtained a valuable understanding of the relationship between complexity metrics and vulnerability detection in Solidity smart contracts. From our findings, we can draw the following conclusions:
@@
    Individual complexity metrics alone are insufficient to predict vulnerable contracts effectively, as their correlations with the existence of vulnerabilities are weak.
@@
This study provides practical evidence of the effectiveness and relevance of complexity metrics in predicting vulnerabilities in Solidity smart contracts and analyses the power of each metric to help researchers understand their individual influence. Furthermore, this study discusses the methods of how complexity metrics can be integrated into existing vulnerability tools as well as how developers can influence these findings and implement a safer code.
@@
The implementations, source code, data, and experimental results are publicly available at https://zenodo.org/records/13364187 and https://github.com/MasoudJTehrani/SC-Complexity-Vulnerability

### v4

@@ Abstract
Software built on poor structural patterns often shows higher exposure to security defects. When code differs from established best practices, verification and maintenance become increasingly difficult, thereby raising the risk of unintentional vulnerabilities. In the context of blockchain technology, where immutable smart contracts handle high-value transactions, the need for strict security assurance is important. This research analyzes the utility of software complexity metrics as diagnostic tools for identifying vulnerable Solidity smart contracts. We evaluate the hypothesis that complexity measures serve as vital, complementary signals for security assessment. Through an empirical examination of 21 distinct metrics, we analyzed their inter-dependencies, statistical association with vulnerabilities, and discriminative capabilities. Our findings indicate a significant degree of redundancy among certain metrics and a relatively low correlation between any single metric and the presence of vulnerabilities. However, the data demonstrates that these metrics possess strong power to distinguish between secure and vulnerable code when analyzed collectively. Notably, with only three exceptions, vulnerable contracts consistently exhibited higher mean complexity scores than their neutral counterparts. While our results show a statistical association, we emphasize that complexity is an indicator rather than a direct cause of vulnerability.
@@ I Introduction & Motivation
Objectivity and Quantifiability: Complexity metrics provide an objective and reproducible way to measure code characteristics [14]. Unlike qualitative assessments (e.g., code readability or subjective security evaluations), these metrics can be systematically computed, analyzed, and compared across large datasets. This makes them a suitable choice for large-scale vulnerability detection efforts.
Practical Integration into Automated Tools: Automated vulnerability detection tools rely on measurable features to assess risk [21]. Complexity metrics are easy to extract from code and can be directly integrated into static analysis tools.
@@ II Contribution & Research Questions
  • RQ4: How different are the metrics in vulnerable codes compared to neutral ones? We examine the strength of the association between high complexity values and the presence of vulnerabilities to determine their utility as predictive indicators. This approach enables a more targeted and efficient allocation of resources toward securing smart contracts.
@@ III-C Smart Contract Architecture
smart contract Loss(\$)
UwU Lend 19.3M
Indodax 22M
BinX Exchange 26M
Penpie 27M
Hedgey Finance 44M
BtcTurk Hot Wallet 55M
DMM Bitcoin Private Key 125M
Orbit Chain Bridge 197M
WazirX 230M
PlayDapp 290M
Total 1.03B
@@ VI Experimental Setup
Mann-Whitney U Test: Since software complexity metrics often follow a non-normal distribution, we employed the Mann-Whitney U test [29] (also known as the Wilcoxon rank-sum test). This non-parametric test allows us to assess whether the distribution of metric values in vulnerable contracts differs significantly from that of neutral contracts, without assuming a normal distribution or equal sample sizes.
In this test, if the p-value is smaller than the chosen significance level (in our case 0.05), you reject the null hypothesis and conclude that there is a significant difference between the distributions of the two independent groups [34].
To quantify the magnitude of the differences between vulnerable and neutral contracts, we calculated Cliff’s Delta ($\delta$[33]. This non-parametric effect size measure estimates the probability that a value from the vulnerable group is strictly greater than a value from the neutral group, minus the reverse probability. We interpret the effect sizes using the thresholds defined by Romano et al. [37]: negligible ($|\delta|<0.147$), small ($0.147\leq|\delta|<0.33$), medium ($0.33\leq|\delta|<0.474$), and large ($|\delta|\geq 0.474$).
@@
TABLE IV: Correlations between complexity metrics and the existence of Vulnerabilities
@@
Metrics SC Correlation TP Correlation
@@
TABLE VI: Statistical comparison (Mann-Whitney U) and Effect Size (Cliff’s Delta) between vulnerable and neutral codes
Metrics Discriminative P-Value Cliff’s Delta Effect Size
SLOC yes <0.05 0.7069 Large
LLOC yes <0.05 0.6137 Large
CLOC yes <0.05 -0.2827 Small
NF yes <0.05 0.5676 Large
WMC yes <0.05 0.5979 Large
NL yes <0.05 0.5809 Large
NLE yes <0.05 0.5699 Large
NUMPAR yes <0.05 0.3782 Medium
NOS yes <0.05 0.7253 Large
DIT yes <0.05 0.4118 Medium
NOA yes <0.05 0.4254 Medium
NOD yes <0.05 -0.1807 Small
CBO yes <0.05 0.5432 Large
NA yes <0.05 0.7391 Large
NOI yes <0.05 0.7418 Large
Avg. McCC yes <0.05 0.5324 Large
Avg. NL yes <0.05 0.4833 Large
Avg. NLE yes <0.05 0.4616 Medium
Avg. NUMPAR yes <0.05 -0.1726 Small
Avg. NOS yes <0.05 0.6541 Large
Avg. NOI yes <0.05 0.6492 Large
VII Experimental Results
In this results section, we give answers to each research question asked. For a full report, you can look at our implementation’s source code which is publicly available at GitHub [^4] https://github.com/MasoudJTehrani/SC-Complexity-Vulnerability
  • NOI and Avg. NOI: These are equal due to the limited use of invocations in smart contracts, suggesting that only one metric might be necessary. The high correlation between NOI and other metrics of SLOC, NF, WMC, and NOS was mentioned previously.
RQ3: Are the metrics different in vulnerable and neutral codes? We sought to investigate whether the complexity metrics can effectively distinguish between functions with reported vulnerabilities and those without reported vulnerabilities. To assess whether the metrics in each group (vulnerable and neutral) can effectively discriminate between vulnerable and neutral codes, we employed the Mann-Whitney U test present in Table VI. This statistical test was conducted by comparing the metric values of the vulnerable group with their corresponding values in the neutral group. While the Mann-Whitney U test indicated statistically significant differences for all metrics ($p<0.05$), confirming their ability to distinguish between vulnerable and neutral codes, the effect size analysis reveals the strength of these distinctions. Notably, no metric exhibited a negligible effect size, and the majority of complexity metrics demonstrated a large effect size. This reinforces the findings of RQ3, suggesting that most complexity measures are not just statistically significant but are practically robust indicators for identifying vulnerable contracts. Metrics with medium effect sizes, such as DIT ($\delta=0.4118$), NOA ($\delta=0.4254$), and Avg. NLE ($\delta=0.4616$), further support the hypothesis that inheritance depth and nesting levels are consistent discriminators, even if slightly less dominant than the primary complexity indicators.
@@ VII Experimental Results
To answer this question, we examine the central tendency and spread of each group. To compare the means of each group, we employed confidence intervals, ensuring that the resulting intervals encompass the true population parameter in around 95% of cases. This statistical approach allowed us to effectively evaluate the differences between vulnerable and neutral codes in Solidity smart contracts. The comparison results are presented in Figure 5.
@@
This finding is also supported by Table VI, with the effect sizes that highlight distinct behaviors among specific features. While most metrics show a strong positive association with vulnerability, CLOC ($\delta=-0.2827$), NOD ($\delta=-0.1807$), and Avg. NUMPAR ($\delta=-0.1726$), exhibited small effect sizes with negative Cliff’s Delta values. This confirms the ”inverse” relationship noted in our conclusions: unlike general complexity, vulnerable contracts tend to have fewer comment lines (CLOC) and fewer descendants (NOD) than their neutral counterparts. This nuance is critical for building predictive models, as it suggests that while high complexity generally signals risk, a lack of documentation (comments) or extensibility (descendants) is also a specific, albeit weaker, marker of vulnerability.
@@ VIII-C Future Work
In future work, we have identified several areas for improvement and expansion in our study. One of the objectives is to include additional complexity metrics that are specific to smart contracts, such as the amount of gas expended, to enhance our analysis.
@@
Furthermore, it’s reasonable to explore the inclusion of comprehension complexity metrics, such as cognitive complexity [11, 12], in our evaluation. Cognitive complexity measures the difficulty of understanding code by considering factors such as nested conditions, control flow, and human cognition. By including such metrics, we can gain a better understanding of the cognitive load imposed on developers and its application to vulnerabilities of Smart Contracts. These future efforts will contribute to a more comprehensive and robust assessment of Smart Contract security.
Future work should also categorize smart contracts by type (e.g., DeFi, Games, Wallets). A contract with high SLOC in a simple domain might be a better predictor of vulnerability than high SLOC in a complex domain.
To move beyond correlation and establish causality, we propose employing longitudinal analyses of code evolution (tracking metrics as vulnerabilities are introduced or fixed) or quasi-experimental designs where complexity is controlled while monitoring security outcomes.
@@ IX-B Internal Threats
This paper employed statistical hypothesis tests, including Mann-Whitney U, to identify significant differences between independent data sets. Since these tests rely on several assumptions, their reliability, particularly for RQ3, may be affected if those assumptions are not met in a different setup. To mitigate this, these methods were selected based on prior research and the characteristics of the dataset. This also ensured consistency with studies used for comparison.
@@
We assessed code complexity metrics (e.g., SLOC, WMC) but did not normalize these against the inherent difficulty of the task the smart contract solves. For example, a complex DeFi protocol naturally requires more lines of code and coupling than a simple token contract. Consequently, higher complexity metrics might simply reflect a larger problem domain, which statistically increases the surface area for bugs, rather than indicating that the coding style itself was poor. While we demonstrate a strong association between these metrics and vulnerabilities, future studies should stratify contracts by domain (e.g., DeFi, NFT, Gaming) or normalize metrics against gas consumption profiles to isolate unnecessary implementation complexity from inherent problem complexity.
@@ X Conclusion
Throughout our investigation, we have obtained a valuable understanding of the relationship between complexity metrics and vulnerability detection in Solidity smart contracts.
It is important to note that these findings represent correlations. High complexity often accompanies larger ‘problem complexity,’ which may be the underlying driver for increased vulnerability risk, rather than the code complexity metrics themselves acting as a causal mechanism. From our findings, we can draw the following conclusions:
@@
    Individual complexity metrics alone are insufficient to help predict vulnerable contracts effectively, as their correlations with the existence of vulnerabilities are weak.
@@
This study provides practical evidence of the effectiveness and relevance of complexity metrics in predicting vulnerabilities in Solidity smart contracts and analyzes the power of each metric to help researchers understand their individual influence. Furthermore, this study discusses the methods of how complexity metrics can be integrated into existing vulnerability tools, as well as how developers can influence these findings and implement safer code.
@@
The implementations, source code, data, and experimental results are publicly available at https://github.com/MasoudJTehrani/SC-Complexity-Vulnerability